# OpenMMReasoner: Pushing the Frontiers for Multimodal Reasoning with an Open and General Recipe

### v1

@@ 1 Introduction
With the rapid progress of reinforcement learning with verifiable rewards (RLVR), leading models such as DeepSeek-R1 [dsr1_cite] and OpenAI o3 [openai_o3_o4mini_2025] have demonstrated strong reasoning capabilities across mathematics, programming, and scientific tasks. These advancements highlight RLVR as an effective paradigm for improving structured reasoning in large language models (LLMs). Motivated by this progress, researchers have increasingly explored RL-enhanced reasoning from LLMs to large multimodal models (LMMs). Recent studies [meng2025mm, chen2025sftrlearlyinvestigation, deng2025openvlthinkercomplexvisionlanguagereasoning, wei2025openvisionreasonertransferring] push the boundaries of multimodal reasoning, showing that reinforcement learning (RL) can strengthen fine-grained visual understanding and complex cross-modal problem-solving. These developments suggest that the benefits of RLVR can transfer beyond text, offering a scalable path toward more capable and reliable multimodal reasoning systems.
Despite these advancements, transparency across the training pipeline remains limited. While numerous studies investigate the supervised-finetuning (SFT) and RL stages [bai2025qwen25vltechnicalreport, openai_o3_o4mini_2025, comanici2025gemini], few provide details of their data curation processes or conduct comprehensive ablation analyses. This lack of openness restricts reproducibility and obscures a deeper understanding of how reasoning-capable LMMs are actually built and how their training dynamics evolve. Recent efforts such as ThinkLite-VL [wang2025thinklite] and MM-Eureka [meng2025mm] propose clearer methodologies for RL data construction and offer useful insights for algorithm design. Other work, including Open Vision Reasoner (OVR) [wei2025openvisionreasonertransferring], attempts to address both SFT and RL, but does not provide a scalable, unified recipe that generalizes across tasks and modalities. Similarly, LLM-focused efforts [guha2025openthoughtsdatarecipesreasoning] emphasize SFT while offering limited discussion of RL or multimodal extensions. Collectively, these limitations point to a critical gap: the lack of a transparent, scalable, and training recipe that unifies SFT and RL for building multimodal reasoning systems.
In this work, we present a comprehensive and scalable empirical study on training Large Multimodal Reasoning Models (LMRMs) from open-source LMMs (e.g., Qwen2.5-VL [bai2025qwen25vltechnicalreport]). Through extensive experimentation, we develop simple yet effective data pipelines for constructing high-quality SFT and RL datasets, enabling reliable cold-start training of a strong LMRM. Our analysis provides detailed insights into which components most effectively contribute at each stage, revealing practical principles for building robust, generalizable, and efficient multimodal reasoning systems. Furthermore, we explore how to maximize scaling efficiency across both SFT and RL phases and summarize key lessons learned throughout the process.
@@
To summarize, our contributions are threefold: (1) Comprehensive Data Curation Insights. We present the first systematic study on curating high-quality SFT and RL data for multimodal reasoning, supported by extensive and rigorous experiments across diverse modalities and reasoning types. We found that scaling data diversity is a critical factor for curating high-quality datasets. While diversity in data sources is important, diversity in answers represents an additional essential axis for improvement. (2) A Strong SFT Recipe for Reasoning. We present a robust and reproducible SFT recipe that effectively equips LMMs with strong reasoning capabilities. Our approach incorporates step-by-step validation and offers practical insights for scaling a high-quality data pipeline focused on reasoning. By carefully selecting an appropriate teacher model for rejection sampling, and incorporating cross-domain data sources, we construct a cold-start dataset that exhibits both high diversity and quality, forming a solid reasoning foundation for subsequent RL. (3) An Advanced RL Recipe for Reasoning Enhancement. We conduct a comprehensive comparative analysis of multiple RL strategies, including GSPO [zheng2025groupsequencepolicyoptimization], GRPO [shao2024deepseekmathpushinglimitsmathematical], and DAPO [yu2025dapoopensourcellmreinforcement], to evaluate their stability, efficiency, and scaling behavior. By selecting the most suitable algorithm, we establish a robust RL pipeline that further sharpens and stabilizes reasoning abilities, delivering both high stability and superior performance.
@@
Qwen2.5-VL-7B-Instruct [bai2025qwen25vltechnicalreport]
M2-Reasoning [an2025llava]
MiMo-VL [coreteam2025mimovltechnicalreport]
OpenVisionReasoner [wei2025openvisionreasonertransferring]
@@ 2 Related Work
RL has emerged as a powerful approach for enhancing reasoning in LLMs. Recent systems such as OpenAI’s o1 [openaio1_cite] and DeepSeek-R1 [dsr1_cite] have shown that multi-step reasoning and self-verification can emerge purely from large-scale RL. Building on these advances, recent works [meng2025mm, deng2025openvlthinkercomplexvisionlanguagereasoning, wei2025openvisionreasonertransferring, chen2025sftrlearlyinvestigation, leng2025mmr1enhancingmultimodalreasoning, qiao2025wemath20versatilemathbook, xu2025llavacotletvisionlanguage, yuan2025vlcogitoprogressivecurriculumreinforcement] extend RL-enhanced reasoning to multimodal models, while SFT studies [xu2025llavacotletvisionlanguage, yang2025r1onevisionadvancinggeneralizedmultimodal, li2025zebracotdatasetinterleavedvision] highlight the importance of high-quality supervision.
Despite this progress, many prior works [wei2025openvisionreasonertransferring, ai2025m2] lack transparent and reproducible training pipelines, and related efforts [guha2025openthoughtsdatarecipesreasoning] focus only on textual reasoning. This limits reproducibility and obscures how data curation and training design affect outcomes. To address this, we present OpenMMReasoner, a scalable and fully open recipe covering both SFT and RL, offering practitioners clear guidance for building reliable multimodal reasoning models.
@@ 3.1 Data Curation
We begin with approximately 103k raw question–answer pairs collected from public datasets including LLaVA-CoT [xu2025llavacotletvisionlanguage], OpenVLThinker [deng2025openvlthinkercomplexvisionlanguagereasoning], and We-Math2.0 [qiao2024wemathdoeslargemultimodal]. These samples mainly cover general VQA and reasoning-oriented tasks. Additional sources used later in the domain mixing stage (Section 3.5) include MMR1 [zhang2024lmmsevalrealitycheckevaluation] for image-based mathematical reasoning and MiroMind-M1 [li2025miromindm1opensourceadvancementmathematical] for text-based mathematical reasoning.
@@
Data sourced from different benchmarks often follow inconsistent answer styles and reasoning formats, which may introduce training instability [ye2025limoreasoning, li2025miromindm1opensourceadvancementmathematical]. We standardize all samples to a unified reasoning format, normalizing textual structure and ensuring consistent step-wise outputs before entering the distillation stage.
@@ 3.2 Experimental Setup
For the SFT training, we adopt online packing and the Liger-Kernel [hsu2025ligerkernel] to accelerate training efficiency. All experiments are conducted using the LMMs-Engine [lmms_engine2025] as the training framework and LMMs-Eval [lmms_eval2024, zhang2024lmmsevalrealitycheckevaluation] for standardized evaluation. We use Qwen2.5-VL-7B-Instruct [bai2025qwen25vltechnicalreport] as our initial checkpoint to start, which also serves as our baseline. Each model is trained until convergence, and further implementation details are provided in the Appendix.
@@ 3.3 Data Distillation
Prior work [dsr1_cite, guha2025openthoughtsdatarecipesreasoning] shows that distillation from a stronger model improves data quality and data efficiency. To identify the most suitable teacher, we compare answer traces distilled from multiple candidates, applying both format validation and answer verification. Only samples whose final answers pass a rule-based validator and an LLM-as-judge check are retained, producing around 59k verified traces. As shown in Tab.˜2, all teacher-based variants outperform the baseline with a clear leading, confirming that a stronger teacher yields higher-quality supervision. Both of the stronger teacher models provide an average gain of at least 4.5 points across all benchmarks, with Qwen3-VL-235B-Instruct [qwen3vl2025] achieving the highest performance and selected as the teacher for distillation. This stage establishes the foundation for larger-scale answer sampling.
@@ 3.4 Length and Difficulty Filtering
OpenAI-GPT-4o [hurst2024gpt] - - 59.9 31.1 - 40.6 34.5 - 64.4 - - - - -
GPT-4o mini [hurst2024gpt] - - 55.1 27.3 - 30.0 31.6 48.8 41.4 - 37.6 - 34.1 74.9
@@
LLaVA-OneVision-7B [li2024llavaonevisioneasyvisualtask] 4.8M - 62.6 17.6 - 17.6 9.0 - 32.0 - 24.1 - 23.6 48.7
InternVL3-8B [zhu2025internvl3exploringadvancedtraining] - - 70.5 28.6 - 33.9 23.0 - 43.6 - - - 37.6 73.6
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] - - 69.2 25.5 25.6 41.1 21.8 53.1 47.9 51.8 37.9 35.1 36.4 67.3
LLaVA-OneVision-1.5-8B [an2025llava] 105M - 69.6 25.6 21.7 46.3 19.8 49.4 45.8 55.4 37.4 25.2 37.0 74.1
@@
VLAA-Thinker-Qwen2.5-7B [chen2025sftrlearlyinvestigation] 126k 25k 68.0 26.4 - 48.2 22.4 - 48.5 - - - - -
ThinkLite-7B-VL [wang2025thinklite] - 11k 71.6 24.6 - 42.9 16.5 - 42.7 - - - - -
VL-Rethinker-7B [wang2025vlrethinkerincentivizingselfreflectionvisionlanguage] - 39k 73.7 28.4 - 46.4 17.8 - 42.7 - 41.7 - - -
M2-Reasoning [ai2025m2] 6.2M 102k 75.0 42.1 - 40.4 - - 50.6 - - - - -
MMR1 [leng2025mmr1enhancingmultimodalreasoning] 1.6M 15k 72.0 31.8 29.0 55.4 27.9 68.0 48.9 52.4 41.1 37.1 43.5 71.1
OpenVLThinker-7B [deng2025openvlthinkercomplexvisionlanguagereasoning] 3.3k 9.6k 65.3 23.0 26.9 38.1 16.8 61.9 44.5 55.1 39.7 38.4 41.0 69.2
MM-Eureka-Qwen-7B [meng2025mm] - 15.6k 72.6 28.1 32.1 45.4 23.0 59.8 46.3 54.4 40.1 37.1 42.4 74.1
OVR-7B [wei2025open] 2M 300k 72.1 51.8 38.2 54.6 33.5 64.8 54.8 51.8 50.2 29.1 44.5 73.6
@@ 3.5 Domain Mixing
To further enhance reasoning generalization, we incorporate supervised data from additional reasoning domains. While the 583k distilled dataset already provides strong multimodal reasoning, its coverage of mathematical reasoning remains limited. We therefore integrate MMR1 [leng2025mmr1enhancingmultimodalreasoning] (image-based math reasoning) and MiroMind-M1 [li2025miromindm1opensourceadvancementmathematical] (text-based math reasoning). As shown in Tab.˜5, adding both types of mathematical supervision consistently improves performance across multimodal and reasoning benchmarks. Combining the general SFT data with these additional mathematical datasets yields our final 874k mixed SFT dataset.
@@ 3.6 Analysis and Insights
In summary, the dataset evolves through three major stages—from 103k raw questions, to 583k distilled general-reasoning samples, and finally to an 874k mixed SFT dataset incorporating both general and mathematical reasoning. As shown in Tab.˜6, our results across nine reasoning benchmarks [qiao2024wemathdoeslargemultimodal, qiao2024wemathdoeslargemultimodal, lu2024mathvistaevaluatingmathematicalreasoning, xiao2024logicvistamultimodalllmlogical, yue2024mmmumassivemultidisciplinemultimodal, yue2025mmmuprorobustmultidisciplinemultimodal, wang2024charxivchartinggapsrealistic, zou2025dynamathdynamicvisualbenchmark] show that our method achieves superior performance and data efficiency compared to other SFT approaches, demonstrating the effectiveness of our recipe in building a strong reasoning base model.
@@ 4.1 Preliminaries of Reinforcement Learning
Group Relative Policy Optimization (GRPO)[shao2024deepseekmathpushinglimitsmathematical] is an efficient variant of PPO that removes the need for a critic network and Generalized Advantage Estimation (GAE), thereby reducing both memory usage and computational overhead. GRPO normalizes rewards within each rollout group to reduce variance and incorporates likelihood-ratio clipping with a KL-divergence penalty to constrain $\pi_{\theta}$ close to the initial SFT policy. The GRPO objective is defined as:
@@
Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) [yu2025dapoopensourcellmreinforcement] addresses several limitations of GRPO, including entropy collapse, training instability, and length bias caused by sample-level loss. To mitigate these issues, DAPO introduces a decoupled clipping mechanism and a dynamic sampling strategy, leading to a more stable and balanced optimization objective:
@@
Group Sequence Policy Optimization (GSPO) [zheng2025groupsequencepolicyoptimization] tackles the token-level importance bias inherent in GRPO by introducing a sequence-level importance ratio for optimization. Additionally, GSPO employs a smaller clipping threshold $\varepsilon$ to enhance training stability. The resulting optimization objective for GSPO is defined as:
@@ 4.2 Dataset Curation
Similar to the SFT stage, we begin by collecting diverse data samples across multiple domains. Our sources include MMEureka [meng2025mm], ViRL [wang2025vlrethinkerincentivizingselfreflectionvisionlanguage], TQA [DBLP:conf/cvpr/KembhaviSSCFH17], We-Math [qiao2025wemath20versatilemathbook], PuzzleVQA [chia2024puzzlevqadiagnosingmultimodalreasoning], AlgoPuzzleVQA [ghosal2024algopuzzlevqa], and ThinkLiteVL [wang2025thinklite], covering domains such as science, mathematics, charts, and puzzles.
@@ 4.3 Experimental Setup
For the RL training stage, we utilize verl [sheng2024hybridflow] and vllm [kwon2023efficient] to accelerate the training process. The evaluation protocol remains consistent with that of the SFT stage, employing LMMs-Eval [lmms_eval2024, zhang2024lmmsevalrealitycheckevaluation] to ensure a unified and standardized evaluation setup. By default, we use temperature 1.0 and the top-performing checkpoint from the SFT phase as the starting point for RL, unless stated otherwise. The checkpoint corresponds to the best-performing model reported in Tab.˜5, where cross-domain mixing of image and text-based math data were applied, along with $\times 8$ sampling for answer traces.

### v2

@@ 1 Introduction
With the rapid progress of reinforcement learning with verifiable rewards (RLVR), leading models such as DeepSeek-R1 [10] and OpenAI o3 [25] have demonstrated strong reasoning capabilities across mathematics, programming, and scientific tasks. These advancements highlight RLVR as an effective paradigm for improving structured reasoning in large language models (LLMs). Motivated by this progress, researchers have increasingly explored RL-enhanced reasoning from LLMs to large multimodal models (LMMs). Recent studies [24, 4, 7, 37] push the boundaries of multimodal reasoning, showing that reinforcement learning (RL) can strengthen fine-grained visual understanding and complex cross-modal problem-solving. These developments suggest that the benefits of RLVR can transfer beyond text, offering a scalable path toward more capable and reliable multimodal reasoning systems.
Despite these advancements, transparency across the training pipeline remains limited. While numerous studies investigate the supervised-finetuning (SFT) and RL stages [3, 25, 6], few provide details of their data curation processes or conduct comprehensive ablation analyses. This lack of openness restricts reproducibility and obscures a deeper understanding of how reasoning-capable LMMs are actually built and how their training dynamics evolve. Recent efforts such as ThinkLite-VL [35] and MM-Eureka [24] propose clearer methodologies for RL data construction and offer useful insights for algorithm design. Other work, including Open Vision Reasoner (OVR) [37], attempts to address both SFT and RL, but does not provide a scalable, unified recipe that generalizes across tasks and modalities. Similarly, LLM-focused efforts [9] emphasize SFT while offering limited discussion of RL or multimodal extensions. Collectively, these limitations point to a critical gap: the lack of a transparent, scalable, and training recipe that unifies SFT and RL for building multimodal reasoning systems.
In this work, we present a comprehensive and scalable empirical study on training Large Multimodal Reasoning Models (LMRMs) from open-source LMMs (e.g., Qwen2.5-VL [3]). Through extensive experimentation, we develop simple yet effective data pipelines for constructing high-quality SFT and RL datasets, enabling reliable cold-start training of a strong LMRM. Our analysis provides detailed insights into which components most effectively contribute at each stage, revealing practical principles for building robust, generalizable, and efficient multimodal reasoning systems. Furthermore, we explore how to maximize scaling efficiency across both SFT and RL phases and summarize key lessons learned throughout the process.
@@
To summarize, our contributions are threefold: (1) Comprehensive Data Curation Insights. We present the first systematic study on curating high-quality SFT and RL data for multimodal reasoning, supported by extensive and rigorous experiments across diverse modalities and reasoning types. We found that scaling data diversity is a critical factor for curating high-quality datasets. While diversity in data sources is important, diversity in answers represents an additional essential axis for improvement. (2) A Strong SFT Recipe for Reasoning. We present a robust and reproducible SFT recipe that effectively equips LMMs with strong reasoning capabilities. Our approach incorporates step-by-step validation and offers practical insights for scaling a high-quality data pipeline focused on reasoning. By carefully selecting an appropriate teacher model for rejection sampling, and incorporating cross-domain data sources, we construct a cold-start dataset that exhibits both high diversity and quality, forming a solid reasoning foundation for subsequent RL. (3) An Advanced RL Recipe for Reasoning Enhancement. We conduct a comprehensive comparative analysis of multiple RL strategies, including GSPO [50], GRPO [30], and DAPO [44], to evaluate their stability, efficiency, and scaling behavior. By selecting the most suitable algorithm, we establish a robust RL pipeline that further sharpens and stabilizes reasoning abilities, delivering both high stability and superior performance.
@@
Qwen2.5-VL-7B-Instruct [3]
M2-Reasoning [2]
MiMo-VL [32]
OpenVisionReasoner [37]
@@ 2 Related Work
RL has emerged as a powerful approach for enhancing reasoning in LLMs. Recent systems such as OpenAI’s o1 [26] and DeepSeek-R1 [10] have shown that multi-step reasoning and self-verification can emerge purely from large-scale RL. Building on these advances, recent works [24, 7, 37, 4, 17, 28, 40, 45, 42] extend RL-enhanced reasoning to multimodal models, while SFT studies [40, 41, 18] highlight the importance of high-quality supervision.
Despite this progress, many prior works [37, 1] lack transparent and reproducible training pipelines, and related efforts [9] focus only on textual reasoning. This limits reproducibility and obscures how data curation and training design affect outcomes. To address this, we present OpenMMReasoner, a scalable and fully open recipe covering both SFT and RL, offering practitioners clear guidance for building reliable multimodal reasoning models.
@@ 3.1 Data Curation
We begin with approximately 103k raw question–answer pairs collected from public datasets including LLaVA-CoT [40], OpenVLThinker [7], and We-Math2.0 [27]. These samples mainly cover general VQA and reasoning-oriented tasks. Additional sources used later in the domain mixing stage (Section 3.5) include MMR1 [48] for image-based mathematical reasoning and MiroMind-M1 [21] for text-based mathematical reasoning.
@@
Data sourced from different benchmarks often follow inconsistent answer styles and reasoning formats, which may introduce training instability [43, 21]. We standardize all samples to a unified reasoning format, normalizing textual structure and ensuring consistent step-wise outputs before entering the distillation stage.
@@ 3.2 Experimental Setup
For the SFT training, we adopt online packing and the Liger-Kernel [11] to accelerate training efficiency. All experiments are conducted using the LMMs-Engine [22] as the training framework and LMMs-Eval [19, 48] for standardized evaluation. We use Qwen2.5-VL-7B-Instruct [3] as our initial checkpoint to start, which also serves as our baseline. Each model is trained until convergence, and further implementation details are provided in the Appendix.
@@ 3.3 Data Distillation
Prior work [10, 9] shows that distillation from a stronger model improves data quality and data efficiency. To identify the most suitable teacher, we compare answer traces distilled from multiple candidates, applying both format validation and answer verification. Only samples whose final answers pass a rule-based validator and an LLM-as-judge check are retained, producing around 59k verified traces. As shown in Tab.˜2, all teacher-based variants outperform the baseline with a clear leading, confirming that a stronger teacher yields higher-quality supervision. Both of the stronger teacher models provide an average gain of at least 4.5 points across all benchmarks, with Qwen3-VL-235B-Instruct [29] achieving the highest performance and selected as the teacher for distillation. This stage establishes the foundation for larger-scale answer sampling.
@@ 3.4 Length and Difficulty Filtering
OpenAI-GPT-4o [14] - - 59.9 31.1 - 40.6 34.5 - 64.4 - - - - -
GPT-4o mini [14] - - 55.1 27.3 - 30.0 31.6 48.8 41.4 - 37.6 - 34.1 74.9
@@
LLaVA-OneVision-7B [20] 4.8M - 62.6 17.6 - 17.6 9.0 - 32.0 - 24.1 - 23.6 48.7
InternVL3-8B [51] - - 70.5 28.6 - 33.9 23.0 - 43.6 - - - 37.6 73.6
Qwen2.5-VL-7B [3] - - 69.2 25.5 25.6 41.1 21.8 53.1 47.9 51.8 37.9 35.1 36.4 67.3
LLaVA-OneVision-1.5-8B [2] 105M - 69.6 25.6 21.7 46.3 19.8 49.4 45.8 55.4 37.4 25.2 37.0 74.1
@@
VLAA-Thinker-Qwen2.5-7B [4] 126k 25k 68.0 26.4 - 48.2 22.4 - 48.5 - - - - -
ThinkLite-7B-VL [35] - 11k 71.6 24.6 - 42.9 16.5 - 42.7 - - - - -
VL-Rethinker-7B [34] - 39k 73.7 28.4 - 46.4 17.8 - 42.7 - 41.7 - - -
M2-Reasoning [1] 6.2M 102k 75.0 42.1 - 40.4 - - 50.6 - - - - -
MMR1 [17] 1.6M 15k 72.0 31.8 29.0 55.4 27.9 68.0 48.9 52.4 41.1 37.1 43.5 71.1
OpenVLThinker-7B [7] 3.3k 9.6k 65.3 23.0 26.9 38.1 16.8 61.9 44.5 55.1 39.7 38.4 41.0 69.2
MM-Eureka-Qwen-7B [24] - 15.6k 72.6 28.1 32.1 45.4 23.0 59.8 46.3 54.4 40.1 37.1 42.4 74.1
OVR-7B [38] 2M 300k 72.1 51.8 38.2 54.6 33.5 64.8 54.8 51.8 50.2 29.1 44.5 73.6
@@ 3.5 Domain Mixing
To further enhance reasoning generalization, we incorporate supervised data from additional reasoning domains. While the 583k distilled dataset already provides strong multimodal reasoning, its coverage of mathematical reasoning remains limited. We therefore integrate MMR1 [17] (image-based math reasoning) and MiroMind-M1 [21] (text-based math reasoning). As shown in Tab.˜5, adding both types of mathematical supervision consistently improves performance across multimodal and reasoning benchmarks. Combining the general SFT data with these additional mathematical datasets yields our final 874k mixed SFT dataset.
@@ 3.6 Analysis and Insights
In summary, the dataset evolves through three major stages—from 103k raw questions, to 583k distilled general-reasoning samples, and finally to an 874k mixed SFT dataset incorporating both general and mathematical reasoning. As shown in Tab.˜6, our results across nine reasoning benchmarks [27, 27, 23, 39, 46, 47, 36, 52] show that our method achieves superior performance and data efficiency compared to other SFT approaches, demonstrating the effectiveness of our recipe in building a strong reasoning base model.
@@ 4.1 Preliminaries of Reinforcement Learning
Group Relative Policy Optimization (GRPO)[30] is an efficient variant of PPO that removes the need for a critic network and Generalized Advantage Estimation (GAE), thereby reducing both memory usage and computational overhead. GRPO normalizes rewards within each rollout group to reduce variance and incorporates likelihood-ratio clipping with a KL-divergence penalty to constrain $\pi_{\theta}$ close to the initial SFT policy. The GRPO objective is defined as:
@@
Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) [44] addresses several limitations of GRPO, including entropy collapse, training instability, and length bias caused by sample-level loss. To mitigate these issues, DAPO introduces a decoupled clipping mechanism and a dynamic sampling strategy, leading to a more stable and balanced optimization objective:
@@
Group Sequence Policy Optimization (GSPO) [50] tackles the token-level importance bias inherent in GRPO by introducing a sequence-level importance ratio for optimization. Additionally, GSPO employs a smaller clipping threshold $\varepsilon$ to enhance training stability. The resulting optimization objective for GSPO is defined as:
@@ 4.2 Dataset Curation
Similar to the SFT stage, we begin by collecting diverse data samples across multiple domains. Our sources include MMEureka [24], ViRL [34], TQA [15], We-Math [28], PuzzleVQA [5], AlgoPuzzleVQA [8], and ThinkLiteVL [35], covering domains such as science, mathematics, charts, and puzzles.
@@ 4.3 Experimental Setup
For the RL training stage, we utilize verl [31] and vllm [16] to accelerate the training process. The evaluation protocol remains consistent with that of the SFT stage, employing LMMs-Eval [19, 48] to ensure a unified and standardized evaluation setup. By default, we use temperature 1.0 and the top-performing checkpoint from the SFT phase as the starting point for RL, unless stated otherwise. The checkpoint corresponds to the best-performing model reported in Tab.˜5, where cross-domain mixing of image and text-based math data were applied, along with $\times 8$ sampling for answer traces.